# Recommender systems, stigmergy, and the tyranny of popularity

### v1

@@ Abstract
Scientific recommender systems, such as Google Scholar and Web of Science, are essential tools for discovery. Search algorithms that power work through stigmergy, a collective intelligence mechanism that surfaces useful paths through repeated engagement. While generally effective, this “rich-get-richer” dynamic results in a small number of high-profile papers that dominate visibility. This essay argues argue that these algorithm over-reliance on popularity fosters intellectual homogeneity and exacerbates structural inequities, stifling innovative and diverse perspectives critical for scientific progress. We propose an overhaul of search platforms to incorporate user-specific calibration, allowing researchers to manually adjust the weights of factors like popularity, recency, and relevance. We also advise platform developers on how word embeddings and LLMs could be implemented in ways that increase user autonomy. While our suggestions are particularly pertinent to aligning recommender systems with scientific values, these ideas are broadly applicable to information access systems in general. Designing platforms that increase user autonomy is an important step toward more robust and dynamic information ecosystems.
@@ Introduction
The model’s role is to characterize search results, respond with potential metadata parameter adjustments and semantic embeddings, then wait for human judgment. Topical fit is still ultimately determined by the scholar reading abstracts and articles; the LLM merely accelerates iteration.

### v2

@@ Abstract
Scientific recommender systems, such as Google Scholar and Web of Science, are essential tools for discovery. Search algorithms that power work through stigmergy, a collective intelligence mechanism that surfaces useful paths through repeated engagement. While generally effective, this “rich-get-richer” dynamic results in a small number of high-profile papers that dominate visibility. This essay argues argue that these algorithm’s over-reliance on popularity fosters intellectual homogeneity and exacerbates structural inequities, stifling innovative and diverse perspectives critical for scientific progress. We propose an overhaul of search platforms to incorporate user-specific calibration, allowing researchers to manually adjust the weights of factors like popularity, recency, and relevance. We also advise platform developers on how text embeddings and LLMs could be implemented in ways that increase user autonomy. While our suggestions are particularly pertinent to aligning recommender systems with scientific values, these ideas are broadly applicable to information access systems in general. Designing platforms that increase user autonomy is an important step toward more robust and dynamic information ecosystems.
@@ Introduction
First, semantic retrieval beyond keywords
Transformer-powered document embeddings capture conceptual similarity that keyword match misses. These embeddings let a query on “stigmergic feedback in scholarly communication” surface adjacent work in evolutionary biology or computer-supported cooperative work—papers whose titles may share no tokens with the query. Empirical benchmarks show that such semantic retrieval widens the candidate set without sacrificing precision Metzler et al., (2021). This enlarges the exploration frontier by reducing reliance on direct keyword matches, instead capturing the gestalt of a query.
Second, LLM-assisted search parameter control
While explicit sliders may feel unnatural, a researcher could direct an LLM agent to adjust the parameter weights according to their preferences. Multi-sentence natural language queries could include both topical directives and requests for ”less popular” or “very recent” publications. Although mediated by an LLM, the search engine would return direct links to indexed documents, so provenance is accessible and audit-ready, and hallucinations should be minimal or non-existent.
@@
The model’s role is to characterize search results, and suggest parameter adjustments and frames for semantic search, then wait for human judgment. Topical fit is still ultimately determined by the scholar reading abstracts and articles; the LLM merely accelerates iteration.